Find the coefficient of x^5 in (1+x)^11.
1
Expert's answer
2020-11-03T18:22:10-0500

"\text{According to the binomial theorem}"

"(1+x)^n= \displaystyle\sum_{k=0}^{n} \binom{n}{k}x^k"

"\text{coefficient of }x^5 \ \binom{11}{5}=\frac{11!}{5!6!}=462"

Answer: 462